15.44 Obtain the convolution of the pairs of signals in Fig. 15.38.

f (t) a f2(t) 0 1 t 0 1 2 4 5 t ILI

can you please flip the rectangle and use that to move through the triangle
Answer #1

= 1 (4) – YE-1) - (1) 8f2C) لم °سمی 1 - (1) . fz (1-2) - - ه ا -4 - 2- ا Cae عاما 1 t2 ج 30 -4 0= (t) پا د او د . 4-5 4-4 0 1Case 2: 047-2 <1 → 2 < t < 3 yolu s rdzy 2- 423 +-5 0 t-21 case 3:- t-520 +-2 >1 2005 t > 3 + was j7a2 = 4-5 o 1 1-2 case 4:-- graph looks like ܂ (8% ܢ 1 + 0:6 - - .... . 1 : 1 2 3 5 5 6 CScanned with CamScanner